# Model order reduction via Lie groups

### v1

@@ 5.4 Optimization Strategy
We investigate a decoupled optimization strategy, in which a first optimization identifies $(G^{*},\Phi^{*})\in\Xi$, and a second optimization identifies the minimizer $\rho^{*}_{\theta,\mu}$ of (47), (50) or (52) for fixed $(G^{*},\Phi^{*})$.
@@ 5.5 Subalgebra search
Theorem 5.6 and (56) allow us to identify a lower dimensional Lie group $H^{*}\subseteq G^{*}$, but this requires that $S_{\mathfrak{g}}\subseteq\mathfrak{h}$ holds for some subalgebra $\mathfrak{h}\subseteq\mathfrak{g}$. In practice, we will relax this condition: we introduce an inner product on $\mathfrak{g}$, and aim to identify a subalgebra $\mathfrak{h}\subseteq\mathfrak{g}$ by principal component analysis (PCA) of $S_{\mathfrak{g}}$ and subsequent completion of a Lie algebra $\mathfrak{h}$ by repeated application of the Lie bracket, such that “most” of $S_{\mathfrak{g}}$ is contained within $\mathfrak{h}$.
@@
6 Examples
Here we provide basic and advanced numerical and analytic examples of the Theory in Sections 4 and 5. For a summary of the numerical examples see Table 1. The code for the numerical examples is available at github.com/YPWotte/MORLie.
Name Section Variables
Radial Oscillator 6.1 $\rho_{\mu}$
Rigid pointcloud 6.2 $G,\rho_{\mu}$
Sheering pointclouds 6.3 $G,\Phi,\rho_{\mu}$
Liver respiration 6.4 $\rho_{\mu}$
Table 1: Summary of degrees of freedom for optimization in numerical examples Sec. 6.1 to 6.4.
6.1 Radial oscillations
We rewrite this as the FOM vector field
$\Phi\big(\alpha,(q_{1},q_{2}))=(q_{1},q_{2}+\alpha)\,$, (59)
where the corresponding infinitesimal generator is
The resulting approximate solutions are
which are compared to FOM trajectories in Figure 2.
6.2 Rigid pointcloud
We consider $H(t)\in SE(3)$ with dynamics
$\dot{H}(t)=H(t)\widetilde{T}(t)\,,\;H(0)=H_{0}\,.$ (65)
for a time-dependent $\widetilde{T}(t)\in se(3)$.
$P_{j,k}=\begin{pmatrix}p^{1}_{j,k}\\ \vdots\\ p^{N}_{j,k}\end{pmatrix}\,$, (67)
and denote the solution snapshots as $S\in\mathbb{R}^{(3N_{i}+1)\times N_{j}N_{k}}$ given by
$g\cdot p^{i}_{j,k}=Ap^{i}_{j,k}+b\,.$ (70)
For fixed $(\textnormal{Aff}(3),\Phi)$, we solve for $\rho^{*}:\mathbb{R}^{3N_{i}}\rightarrow\textnormal{aff}(3)$ in two ways: first by directly applying Theorem 5.4 (in code we use a Lie-group version of the Levenberg-Marquardt algorithm for optimization, see e.g. [61, Chapter 8.4.2]), and second by approximating the FOM vector field as
$X_{\mu}(P_{j,k},t)\approx(P_{j,k}-P_{j+1,k})/\Delta t\,$, (71)
This results in a ROM on the Lie group $SE(3)$:
$\dot{H}(t)=\rho_{\theta^{*}}(t)H(t)\,,\;H(0)=I_{4}\,$, (73)
with reconstructed solutions
$P_{i}(t)=\Phi(H(t),P_{i,0})\,.$ (74)
Snapshots of a sample trajectory of a point-cloud are shown in Figures 3(a) - 3(c), along with the reconstructed trajectory with $S_{\textnormal{aff}(3)}$ based on Theorem 5.4. Figures 4(a) and 4(b) shows the error of the reconstructions via MORLie, Figure 4(c) shows the singular values of $S$ in (68) that relate to the approximation error of a naive POD method and the Kolmogorov $N$-width of the problem, and Figure 4(d) shows the singular values of $S_{\textnormal{aff}(3)}$ that relate to the Kolmogorov $\Xi^{\textnormal{Aff}(3)}_{N}$ width of the problem.
(a) T = 0.01s.
(b) T = 2.5s.
(c) T = 5s.
(c) Singular values of $S$.
(d) Singular values of $S_{\textnormal{aff}(3)}$.
6.3 Sheering pointclouds
$p^{i}_{j,k}=A_{f(j)}(t_{k})p^{i}_{j,0}+b_{f(j)}(t_{k})+\eta^{i}_{j,k}\,.$ (75)
$\dot{g}_{f(j)}=\widetilde{A}_{f(j)}(t)g_{f(j)}\,,\;g_{f(j)}(0)=g_{f(j),0}\,.$ (76)
For an initial optimization inspired by Theorem 5.5, we present Algorithm 2 in Appendix B. This algorithm uses a nearest-neighbor filtering approach to split $S$ into clusters $S_{1},\cdots,S_{n_{G}}$, without further prior knowledge, such that points $p^{i}_{j,k}$ in $S_{f(j)}$ satisfy
$p^{i}_{j,k}\approx A_{f(j)}(t_{k})p^{i}_{j,0}+b_{f(j)}(t_{k})\,.$ (77)
(a) T = 0.01s.
(b) T = 2.5s.
(c) T = 5s.
(c) Singular values of $S$.
(d) Singular values of $S_{\textnormal{aff}^{2}(3)}$.
6.4 Tracking a liver during respiration
(a) T = 0.01s.
(b) T = 58s.
(c) T = 116s.
@@ 6.7 Discussion
The results in Section 6.2 show that a MORLie ROM on a Lie group $SE(3)$ was successfully fit to pointcloud data undergoing a rigid motion, without assuming prior knowledge of the particular Lie group. Figure 4(d) shows that the dimension of the identified Lie group, 6, reflects in the singular values of the reduced snapshot matrix. Instead, Figure 4(c) shows that around 93 terms are required to capture most complexity of the data using POD. However, we note that the twist $\widetilde{T}(t)$ had to be chosen of sufficient complexity for Algorithm 1 to correctly identify the subalgebra $se(3)$. For increasing measurement noise, the Algorithm became less reliable and no longer identified $se(3)$ correctly, instead terminating at $\textnormal{aff}(3)$. Future work may investigate more robust version of Algorithm 1. Similarly, identifying $se(3)$ from Algorithm 1 still required expert knowledge, because the resulting basis elements are non-standard. This may be improved by e.g., comparing the identified sub-algebra against a library of subalgebras.
@@
Both results of Section 6.2 and 6.3 support that the velocity-free optimization, albeit more computationally expensive, leads to more accurate results. The authors explain this by the velocity-free optimization being more robust to the Gaussian noise $\eta$, which is amplified in the velocity-based method through the numerical differentiation (71).
The results in Section 6.4 show that a rigid motion accurately models the livers motion during respiration for shallow breathing (Figures 7(a) and 7(b)), but not for deep breaths (Figure 7(c)), during which the liver significantly deforms. This reflects in the full-trajectory error in Figure 8(a), which is accurate to within 2mm for large parts of the trajectory but increases up to 7mm when deformation plays a large role. Future work may investigate nonlinear group actions to more accurately model deformations during respiration.
Finally, Sections 6.5 and 6.6 highlight that MORLie is applicable to distributed systems, although the presented theory in the main article focused on the finite dimensional case. Here, the key-insight is that fixed basis elements can be realized as actions of $\mathbb{R}^{k}$, and moving basis elements by the action of a product Lie group $G\times\mathbb{R}^{k}$. At this point, the formalism becomes similar to transformation-based methods [33, 34, 35, 36], for which MORLie provides a geometric picture that generalizes to manifolds.
@@ 7 Conclusion
We presented a novel geometric framework for model order reduction via Lie groups (MORLie), giving a perspective on reduced order modeling that diverges from established linear subspace and submanifold methods for MOR. In this picture, the FOM evolves on a differentiable manifold and the ROM is cast on a Lie group acting on the FOM manifold. We defined a generalized notion of Kolmogorov $N$-widths, showed how the familiar notion of Kolmogorov $N$-widths for linear subspaces is recovered as a special case, and how the Kolmogorov $N$-width of the presented methods is generally lower than the linear subspace Kolmogorov $N$-width. We derived optimization methods for intrusive and non-intrusive optimization of the ROM, presented initial ideas for hyperreduction. For the example of rigid and affinely moving point-clouds we computationally implement the MOR process, and show that it is more accurate than a linear subspace approximation, while achieving a significantly lower dimension. Finally, we showed that existing Lie group methods can be recast in the framework of MORLie.
@@ 8 Acknowledgements
We want to thank Ana Cordón Avila for sharing the data on edge-tracking in ultrasound scans of a patient’s liver during respiration.
@@ B.2 Group and action search
Appendix C Model order reduction on manifolds
C.1 ManiMOR

### v2

@@ 5.4 Optimization Strategy
We investigate a decoupled optimization strategy, in which a first optimization identifies $(G^{*},\Phi^{*})\in\Xi$, and a second optimization identifies the minimizer $\rho^{*}_{\theta,\mu}$ of (59), (62) or (64) for fixed $(G^{*},\Phi^{*})$. The following theorem states that if the solution trajectories already lie on a group orbit, then the optimal reduced vector field recovers the FOM exactly.
@@ 5.5 Subalgebra search
Theorem 5.6 and (68) allow us to identify a lower dimensional Lie group $H^{*}\subseteq G^{*}$, but this requires that $S_{\mathfrak{g}}\subseteq\mathfrak{h}$ holds for some subalgebra $\mathfrak{h}\subseteq\mathfrak{g}$.
In practice, however, it rarely holds that $S_{\mathfrak{g}}$ is exactly contained in some subalgebra of $\mathfrak{g}$. We therefore relax the condition and approximate the dominant Lie algebra directions via PCA. Concretely, we introduce an inner product on $\mathfrak{g}$, and aim to identify a subalgebra $\mathfrak{h}\subseteq\mathfrak{g}$ by principal component analysis (PCA) of $S_{\mathfrak{g}}$ and subsequent completion of a Lie algebra $\mathfrak{h}$ by repeated application of the Lie bracket, such that “most” of $S_{\mathfrak{g}}$ is contained within $\mathfrak{h}$.
@@
6 Analytic Examples
Sections 6.1 and 6.2 highlight that MORLie is applicable to distributed systems, although the presented theory in the main article focused on the finite dimensional case.
@@ 6.2 The method of freezing in MORLie
Remark 6.5
Here, the key-insight is that fixed basis elements can be realized as actions of $\mathbb{R}^{k}$, and moving basis elements by the action of a product Lie group $G\times\mathbb{R}^{k}$. At this point, the formalism becomes similar to that of transformation-based methods [33, 34, 35, 36] and the shifted POD [38, 39], for which MORLie provides a geometric picture that generalizes to manifolds.
7 Numerical Examples
Here we provide basic and advanced numerical examples of the Theory in Sections 4 and 5. For a summary see Table 1. The code is available at github.com/YPWotte/MORLie.
Name Section Variables
Radial Oscillator 7.1 $\rho_{\mu}$
Rigid pointcloud 7.2 $G,\rho_{\mu}$
Sheering pointclouds 7.3 $G,\Phi,\rho_{\mu}$
Liver respiration 7.4 $\rho_{\mu}$
Table 1: Summary of degrees of freedom for optimization in numerical examples Sec. 7.1 to 7.4.
7.1 Radial oscillations
We rewrite this as the FOM vector field
$\Phi\big(\alpha,(q_{1},q_{2}))=(q_{1},q_{2}+\alpha)\,$, (104)
where the corresponding infinitesimal generator is
The resulting approximate solutions are
which are compared to FOM trajectories in Figure 2.
7.2 Rigid pointcloud
We consider $H(t)\in SE(3)$ with dynamics
$\dot{H}(t)=H(t)\widetilde{T}(t)\,,\;H(0)=H_{0}\,.$ (110)
for a time-dependent $\widetilde{T}(t)\in se(3)$.
$P_{j,k}=\begin{pmatrix}p^{1}_{j,k}\\ \vdots\\ p^{N}_{j,k}\end{pmatrix}\,$, (112)
and denote the finite set of state-time solution snapshots as $S\in\mathbb{R}^{(3N_{i}+1)\times N_{j}N_{k}}$ given by
7.2.1 Application of a naive POD
For a naive POD, we directly investigate the singular values of $S$ without further pre-processing. Leaving noise aside for an initial analysis:
Theorem 7.1
Assume the following:
  1. 1.
    The number of columns $N_{j}N_{k}$ and the number of rows $3N_{i}+1$ of $S$ are sufficiently large: $N_{j}N_{k}\geq 9(N_{j}+1)+3$ and $3N_{i}+1\geq 9(N_{j}+1)+3$
  2. 2.
    Each initial point cloud $P_{j,0}\in\mathbb{R}^{N_{j}}$ spans $\mathbb{R}^{3}$:
    $\dim\textnormal{span}\{p^{i}_{j,0}\;|\;i\in\{0,\cdots,N_{i}\}\}=3$
  3. 3.
    The matrices $R(t_{k})$ span $\mathbb{R}^{3\times 3}$:
    $\dim\textnormal{span}\{R(t_{k})\;|\;k\in\{0,\cdots,N_{k}\}\}=9$
  4. 4.
    The translations $b(t_{k})$ span $\mathbb{R}^{3}$:
    $\dim\textnormal{span}\{b(t_{k})\;|\;k\in\{0,\cdots,N_{k}\}\}=3$
  5. 5.
    The subspaces
    $W_{j}:=\textnormal{span}\{P_{j,k}\;|\;k\in\{0,\cdots,N_{k}\}\}$
    of $\mathbb{R}^{3N_{i}}$ associated with different point cloud trajectories $P_{j,k}$ are independent.
Then the number of non-zero singular values of the snapshot set $S$ in (113) is $9(N_{j}+1)+3$.
Proof 7.1.1
For the proof, see Appendix C.
Hence, the singular values of $S$ scale with the number of point-cloud trajectories. The assumptions for Theorem 7.1 are not strong: given randomly sampled initial pointclouds $P_{j,0}$ with sufficiently many points, they hold generically. Considering the simplicity of the underlying (6D) group dynamics (110), this is not yet a problem. If the group dynamics were more complicated, and required many different point-cloud trajectories to capture their complexity, the application of a POD without further processing of $S$ would become unfeasible.
Remark 7.2
The noise differs per particle in the point cloud, as is common in practical scenarios, and it is effectively a random vector in $\mathbb{R}^{3N_{i}}$. Expressing the noise accurately requires a POD basis that spans $\mathbb{R}^{3N_{i}}$, leading to a noise floor in the singular values.
7.2.2 Application of MORLie
$g\cdot p^{i}_{j,k}=Ap^{i}_{j,k}+b\,.$ (115)
For fixed $(\textnormal{Aff}(3),\Phi)$, we solve for $\rho^{*}:\mathbb{R}^{3N_{i}}\rightarrow\textnormal{aff}(3)$ in two ways: first by directly applying Theorem 5.4 (in code we use a Lie-group version of the Levenberg-Marquardt algorithm for optimization, see e.g. [66, Chapter 8.4.2]), and second by approximating the FOM vector field as
$X_{\mu}(P_{j,k},t)\approx(P_{j,k}-P_{j+1,k})/\Delta t\,$, (116)
This results in a ROM on the Lie group $SE(3)$:
$\dot{H}(t)=\rho_{\theta^{*}}(t)H(t)\,,\;H(0)=I_{4}\,$, (118)
with reconstructed solutions
$P_{i}(t)=\Phi(H(t),P_{i,0})\,.$ (119)
7.2.3 Results
Snapshots of a sample trajectory of a point-cloud are shown in Figures 3(a) - 3(c), along with the reconstructed trajectory with $S_{\textnormal{aff}(3)}$ based on Theorem 5.4. Figures 4(a) and 4(b) shows the error of the reconstructions via MORLie, and Figure 4(d) shows the error for various projections to a POD basis. Figure 4(c) shows the singular values of $S$ in (113). The singular values sharply drop at $9(N_{j}+1)+3=93$, c.f. Theorem 7.1. The singular values do not drop to zero since noise is added to each node, c.f. Remark 7.2. Figure 4(e) shows the singular values of $S_{\textnormal{aff}(3)}$ that relate to the Kolmogorov $\Xi^{\textnormal{Aff}(3)}_{N}$ width of the problem.
(a) T = 0.01s.
(b) T = 2.5s.
(c) T = 5s.
(c) Singular values of $S$.
(d) Full trajectory errors $\textnormal{dist}\big(P_{POD}(t),P(t)\big)$, where $P_{POD}$ is a projection of $P$ to a POD basis.
(e) Singular values of $S_{\textnormal{aff}(3)}$.
7.3 Sheering pointclouds
$p^{i}_{j,k}=A_{f(j)}(t_{k})p^{i}_{j,0}+b_{f(j)}(t_{k})+\eta^{i}_{j,k}\,.$ (120)
$\dot{g}_{f(j)}=\widetilde{A}_{f(j)}(t)g_{f(j)}\,,\;g_{f(j)}(0)=g_{f(j),0}\,.$ (121)
For an initial optimization inspired by Theorem 5.5, we present Algorithm 2 in Appendix B. This algorithm uses a nearest-neighbor filtering approach to split finite $S$ into clusters $S_{1},\cdots,S_{n_{G}}$, without further prior knowledge, such that points $p^{i}_{j,k}$ in $S_{f(j)}$ satisfy
$p^{i}_{j,k}\approx A_{f(j)}(t_{k})p^{i}_{j,0}+b_{f(j)}(t_{k})\,.$ (122)
7.3.1 Results
(a) T = 0.01s.
(b) T = 2.5s.
(c) T = 5s.
(c) Singular values of $S$.
(d) Singular values of $S_{\textnormal{aff}^{2}(3)}$.
7.4 Tracking a liver during respiration
(a) T = 0.07s.
(b) T = 58s.
(c) T = 116s.
@@ 7.5 Discussion
The results in Section 7.2 show that a MORLie ROM on a Lie group $SE(3)$ was successfully fit to pointcloud data undergoing a rigid motion, without assuming prior knowledge of the particular Lie group. Figure 4(e) shows that the dimension of the identified Lie group, 6, reflects in the singular values of the reduced snapshot matrix. Instead, Section 7.2.1 and Figure 4(c) show that $9(N_{j}+1)+3=93$ terms are required to capture most complexity of the data with a naive POD. Given more complicated underlying group dynamics requiring larger batches of experimental point cloud data, this is expected to make a naive POD unfeasible for data-based ROM identification – increasing the number of experiments should increase ROM accuracy, not the number of terms required.
We further note that the twist $\widetilde{T}(t)$ had to be chosen of sufficient complexity for Algorithm 1 to correctly identify the subalgebra $se(3)$. For increasing measurement noise, the Algorithm became less reliable and no longer identified $se(3)$ correctly, instead terminating at $\textnormal{aff}(3)$. Future work may investigate more robust version of Algorithm 1. Similarly, identifying $se(3)$ from Algorithm 1 still required expert knowledge, because the resulting basis elements are non-standard. This may be improved by e.g., comparing the identified sub-algebra against a library of subalgebras.
@@
Both results of Section 7.2 and 7.3 support that the velocity-free optimization, albeit more computationally expensive, leads to more accurate results. The authors explain this by the velocity-free optimization being more robust to the Gaussian noise $\eta$, which is amplified in the velocity-based method through the numerical differentiation (116).
Finally, the Section 7.4 shows that a rigid motion accurately models the livers motion during respiration for shallow breathing (Figures 7(a) and 7(b)), but not for deep breaths (Figure 7(c)), during which the liver significantly deforms. This reflects in the full-trajectory error in Figure 8(a), which is accurate to within 2mm for large parts of the trajectory but increases up to 7mm when deformation plays a large role. Future work may investigate nonlinear group actions to more accurately model deformations during respiration.
@@ 8 Conclusion
We presented a novel geometric framework for model order reduction via Lie groups (MORLie), giving a perspective on reduced order modeling that diverges from established linear subspace and submanifold methods for MOR. In this picture, the FOM evolves on a differentiable manifold and the ROM is cast on a Lie group acting on the FOM manifold. We defined a generalized notion of Kolmogorov $N$-widths – the group Kolmogorov $N$-width – showed how the familiar notion of Kolmogorov $N$-widths for linear subspaces is recovered as a special case, how the Kolmogorov $N$-width of the presented methods is generally lower than the linear subspace Kolmogorov $N$-width, and related it to recent nonlinear widths. We derived optimization methods for intrusive and non-intrusive optimization of the ROM and presented initial ideas for hyperreduction. In an extensive analytic examples, we apply MORLie to the transport equation and show that existing Lie group methods can be recast in the framework of MORLie. Finally, for the example of rigid and affinely evolving point-clouds we computationally implement the MOR process, and show that it scales better than linear subspace approximations, achieving a consistently low dimension.
@@ 9 Acknowledgements
We want to thank Ana Cordón Avila for sharing the data on edge-tracking in MRI scans of a patient’s liver during respiration.
@@ B.2 Group and action search
Appendix C Singular values of rigid snapshots
We prove Theorem 7.1. For convenience, we state the Theorem again:
Theorem C.1
Assume the following:
  1. 1.
    The number of columns $N_{j}N_{k}$ and the number of rows $3N_{i}+1$ of $S$ are sufficiently large: $N_{j}N_{k}\geq 9(N_{j}+1)+3$ and $3N_{i}+1\geq 9(N_{j}+1)+3$
  2. 2.
    Each initial point cloud $P_{j,0}\in\mathbb{R}^{N_{j}}$ spans $\mathbb{R}^{3}$:
    $\dim\textnormal{span}\{p^{i}_{j,0}\;|\;i\in\{0,\cdots,N_{i}\}\}=3$
  3. 3.
    The matrices $R(t_{k})$ span $\mathbb{R}^{3\times 3}$:
    $\dim\textnormal{span}\{R(t_{k})\;|\;k\in\{0,\cdots,N_{k}\}\}=9$
  4. 4.
    The translations $b(t_{k})$ span $\mathbb{R}^{3}$:
    $\dim\textnormal{span}\{b(t_{k})\;|\;k\in\{0,\cdots,N_{k}\}\}=3$
  5. 5.
    The subspaces
    $W_{j}:=\textnormal{span}\{P_{j,k}\;|\;k\in\{0,\cdots,N_{k}\}\}$
    of $\mathbb{R}^{3N_{i}}$ associated with different point cloud trajectories $P_{j,k}$ are independent.
Then the number of non-zero singular values of the snapshot set $S$ in (113) is $9(N_{j}+1)+3$.
Proof C.1.1
The number of non-zero singular values of a matrix is bounded by the number of columns $N_{j}N_{k}$, the number of rows $3N_{i}+1$, and the column rank of $S$. Given Assumption 1. that $N_{j}N_{k}$ and $3N_{i}+1$ are sufficiently large, we show that the column rank of $S$ is $9(N_{j}+1)+3$.
Each column of $S$ corresponding to trajectory $j$ at time $t_{k}$ has the form
$P_{j,k}=\begin{pmatrix}R(t_{k})p^{1}_{j,0}\\ \vdots\\ R(t_{k})p^{N_{i}}_{j,0}\end{pmatrix}+\begin{pmatrix}b(t_{k})\\ \vdots\\ b(t_{k})\end{pmatrix}$. (144)
We analyze rotational and translational contributions separately, first focusing on the rotational term.
Write as $e_{n}\in\mathbb{R}^{3}$ the $n$-th unit vector, and define vectors $\bar{P}^{(j)}_{nm}\in\mathbb{R}^{3N_{i}}$ by
$\bar{P}^{(j)}_{nm}=\begin{pmatrix}(e_{n}^{\top}p^{1}_{j,0})e_{m}\\ \vdots\\ (e_{n}^{\top}p^{N_{i}}_{j,0})e_{m}\end{pmatrix}\,$. (145)
Then
$P^{\mathrm{rot}}_{j,k}=\begin{pmatrix}R(t_{k})p^{1}_{j,0}\\ \vdots\\ R(t_{k})p^{N_{i}}_{j,0}\end{pmatrix}$ $=\sum_{n,m}R_{nm}(t_{k})\begin{pmatrix}(e_{n}^{\top}p^{1}_{j,0})e_{m}\\ \vdots\\ (e_{n}^{\top}p^{N_{i}}_{j,0})e_{m}\end{pmatrix}$ (146)
$=\sum_{n,m}R_{nm}(t_{k})\bar{P}^{(j)}_{nm}\,$. (147)
Hence the rotational part lies in the span of at most $9$ vectors. We now show that these are linearly independent. Suppose that
$\sum_{n,m=1}^{3}a_{nm}\bar{P}^{(j)}_{nm}=0.$ (148)
Then for each particle $p^{i}_{j,0}$
$\sum_{n,m}a_{nm}(e_{n}^{\top}p^{i}_{j,0})e_{m}=0\,.$ (149)
Define
$A:=\sum_{n,m}a_{nm}e_{m}e_{n}^{\top}\in\mathbb{R}^{3\times 3}.$ (150)
The previous equation is equivalent to
$Ap^{i}_{j,0}=0\quad\text{for all }i.$ (151)
By Assumption 2., the initial cloud spans $\mathbb{R}^{3}$, so this would imply $A=0$, and therefore $a_{nm}=0$ for all $n,m$. Thus, the nine vectors $\bar{P}^{(j)}_{nm}$ are linearly independent.
By Assumption 3., the rotational component for trajectory $j$ spans a $9$-dimensional subspace of $\mathbb{R}^{3N_{i}}$.
By Assumption 5., it follows that $9$ linearly independent components are required for the rotational component of each of the $N_{j}+1$ trajectories, and we arrive at $9(N_{j}+1)$ linearly independent components to express the rotational component of $S$.
By Assumption 4., the translations can be expressed by three linearly independent vectors $\bar{P}_{b,1},\cdots,\bar{P}_{b,3}$ for all trajectories, bringing the maximum number of orthogonal components to $9(N_{j}+1)+3$.
Appendix D Model order reduction on manifolds
D.1 ManiMOR